# On Diversified Preferences of Large Language Model Alignment

### v4

@@ Abstract
Aligning large language models (LLMs) with human preferences has been recognized as the key to improving LLMs’ interaction quality. However, in this pluralistic world, human preferences can be diversified due to annotators’ different tastes, which hinders the effectiveness of LLM alignment methods. This paper presents the first quantitative analysis of commonly used human feedback datasets to investigate the impact of diversified preferences on reward modeling. Our analysis reveals a correlation between the calibration performance of reward models (RMs) and the alignment performance of LLMs. We find that diversified preference data negatively affect the calibration performance of RMs on human-shared preferences, such as Harmless&Helpful, thereby impairing the alignment performance of LLMs. To address the ineffectiveness, we propose a novel Multi-Objective Reward learning method (MORE) to enhance the calibration performance of RMs on shared preferences. We validate our findings by experiments on three models and five human preference datasets. Our method significantly improves the prediction calibration of RMs, leading to better alignment of the Alpaca-7B model with Harmless&Helpful preferences. Furthermore, the connection between reward calibration and preference alignment performance suggests that calibration error can be adopted as a key metric for evaluating RMs. The open-source code and data are available at https://github.com/dunzeng/MORE.
@@
Dun Zeng1,333Work done during an internship at Tencent AI Lab,111Equal contribution. YongDai1,111Equal contribution. Pengyu Cheng1,111Equal contribution. Longyue Wang1 Tianhao Hu1 Wanshun Chen1 Nan Du1 Zenglin Xu2,222Corresponding author: zenglin@gmail.com 1Tencent AI Lab, Shenzhen, China 2Peng Cheng Lab, Shenzhen, China
@@ 1 Introduction
The performance of these alignment methods highly depends on the quality of human preference data , where is the input query to the LLM, and response is preferred to response under the human annotation Ouyang et al. (2022). Ideally, the preference datasets should uniformly be helpful, harmless, benevolent, and unbiased to guide the LLM alignment. However, in real-world scenarios, individuals can have diversified preferences on the same topic based on their different experiences, educational backgrounds, religions, and cultures (Leonardelli et al., 2021). Even for the same person, his or her expected model answer to a particular question can vary depending on different scenarios (Cheng et al., 2023a). The annotation disagreement, which is caused by different annotators or the same annotator in different scenarios Bai et al. (2022), will significantly hinder the effectiveness of alignment methods Davani et al. (2022); Wan et al. (2023).
@@
In our exploration, we found the over-rewarding phenomenon, that is, the vanilla RMs tend to output extreme rewards on samples, which damages the RMs and LLM alignment. To enhance the efficiency of leveraging the diversified preference datasets, inspired by multi-objective optimization methods Sener and Koltun (2018); Zeng et al. (2023), we regard RMs as a shared reward additionally with a customized reward drift. The shared reward represents the shared preferences across datasets (or general human preferences) and the reward drift contains individual or domain-specific preference information (Cheng et al., 2023a). Then, we introduce a Multi-Objective Reward training scheme (MORE) to capture the shared (general) preference information, which adopts a novel reweight techniques to minimize the mean gradient of enlarging reward drifts. With MORE, RMs can capture a broader range of preferences and mitigate the impact of reward drifts. The main contributions of this paper are:
@@ 4.3 Training Scheme: MORE
where we adjust to minimize the partial gradient of enlarging reward drifts. The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023). We provide the details of our efficient implementation in the Appendix B. Furthermore, shares the same magnitude of vanilla loss function in expectation over the whole training dataset, as justified in Appendix B.
@@ 5 Experiments on Reward Modeling
In this section, we present our experiments and quantitative analyses on reward modeling.
@@ 5.1 Reward Modeling on Diversified Preference Datasets
Increasing the size of LLMs, reward model training on mixed diversified preference datasets can maintain reward accuracy. For instance, when Pythia-1.4B is used as the RM base model, the reward accuracy is lower compared to the Top accuracy achieved through single preference training on all preferences. Then, when LLaMa2-7B is used as the base model, the reward accuracy on the Oasst1, Webgpt, and Summarise test sets surpasses the top accuracy achieved through single training. Additionally, the degradation of reward accuracy on the Helpful and Harmless datasets is mitigated. Therefore, the performance of RMs typically is proportional to the size of base models (Gao et al., 2023). Moreover, we find the accuracy of RM is poor, revealing the preference conflicts across RM.
@@
Reward Model Perplexity (PPL) GPT4 Evaluation (%) Base Model Scheme Acc(%) ECE Helpful Harmless Win Tie Lose - - - - 15.48 12.71 - - - Pythia-1.4B MultiTask 64.79 0.0177 15.30 8.22 44 22 34 MORE 64.32 0.0109 12.68 8.42 Pythia-2.8B MultiTask 66.61 0.0145 16.76 8.42 45 21 34 MORE 65.87 0.0078 13.14 10.29 LLaMa2-7B MultiTask 72.40 0.0284 16.93 8.69 45 23 32 MORE 72.32 0.0143 11.97 9.96 Table 1: The RJS alignment performance with different RMs. The first line is the performance of Alpaca base model. The results show that ECE further reflects the ability of RMs when the reward accuracy is close.
@@ 6 Experiments on LLM Alignment
We finetune the Alpaca model on the good response from H&H training datasets, and the finetuned model is marked by Alpaca-SFT. Then, we conduct the RJS alignment experiments with LLaMa2-7B RMs from Figure 6. We compare each alignment result of Alpaca-RJS models with the same Alpaca-STF model via GPT evaluation, shown in Figure 7. The results show that the RMs with lower ECE values work better for RJS alignments, emphasizing the importance of calibration evaluation.
@@ 7 Additional Discussions
This paper reveals two main suggestions for future reward model training works. First, Evaluate RMs with reward accuracy and calibration error. Reward accuracy is insufficient to evaluate the ability of RMs due to model capacity and data quality. Our work suggests the community additionally focuses on the calibration performance of RMs. Besides, Increasing the diversity of preference data samples can ensure the robustness of the reward modeling process. Due to the preference information being typically noisy, learning reward information from mixed diversified datasets can be beneficial.
@@
The MORE can enhance preference modeling pre-trained (PMP) paradigm (Askell et al., 2021) as it captures the shared preference information. This facilitates its use in federated learning scenarios (McMahan et al., 2017), where the data distributions are highly heterogeneous across participants. Moreover, the RM can be easily finetuned to specific preferences (Cheng et al., 2023a). This flexibility allows for the adaptation of our approach to various applications.
@@ Appendix A Related Work
Multi-objective Optimization (MOO) (Gunantara, 2018) is a branch of methods addressing learning problems involving multiple conflicting objectives. In real-world scenarios, it commonly encounters situations where multiple objectives need to be considered simultaneously, often with trade-offs between them. In the practice of machine learning, most MOO methods Sener and Koltun (2018); Zeng et al. (2023) apply linear scalarization Barrett and Narayanan (2008) to merge multiple objectives into one, and then automatically adjust the objective coefficients to balance the conflicts among different tasks.
@@ Appendix B Detailed Discussions about MORE
We use adaptive weighting methods to reduce the reward drift across preferences and adjust the reward modeling process in the data batch-wise. The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023). However, the computing cost of solving it is proportional to the size of parameters . Since the size of is in the billions, we only utilize gradients on the reward head from each preference to avoid expensive computation cost. In detail, we obtain the hidden states before the reward head and compute the gradient of the reward head solely with data . Collecting the reward head gradient from diversified preferences, the is computed by:

### v5

@@ Abstract
Aligning large language models (LLMs) with human preferences has been recognized as the key to improving LLMs’ interaction quality. However, in this pluralistic world, human preferences can be diversified due to annotators’ different tastes, which hinders the effectiveness of LLM alignment methods. This paper presents the first quantitative analysis of the experimental scaling law for reward models with varying sizes, from 1.3 billion to 7 billion parameters, trained with human feedback exhibiting diverse preferences. Our analysis reveals that the impact of diversified human preferences depends on both model size and data size. Larger models with sufficient capacity mitigate the negative effects of diverse preferences, while smaller models struggle to accommodate them. To mitigate the impact of diverse preferences, we introduce a new metric, Expected Calibration Error (ECE), to evaluate RMs and show their obvious positive correlation with the alignment performance of LLMs. Furthermore, we propose a Multi-Objective Reward learning method (MORE) to enhance the calibration performance of RMs on shared preferences. Through experiments on four models and five human preference datasets, we find the calibration error can be adopted as a key metric for evaluating RMs and MORE can obtain superior alignment performance.
@@
Dun Zeng1,111Equal contribution.333Part of work was done at Tencent AI Lab YongDai2,111Equal contribution.333Part of work was done at Tencent AI Lab Pengyu Cheng1,111Equal contribution. Longyue Wang3,333Part of work was done at Tencent AI Lab Tianhao Hu1 Wanshun Chen1 Nan Du1 Zenglin Xu4,222Corresponding author: zenglin@gmail.com 1Tencent AI Lab, 2HiThink Research, 3Alibaba Group, 4Peng Cheng Lab
@@ 1 Introduction
The performance of these alignment methods highly depends on the quality of human preference data , where is the input query to the LLM, and response is preferred to response under the human annotation Ouyang et al. (2022). Ideally, the preference datasets should uniformly be helpful, harmless, benevolent, and unbiased to guide the LLM alignment. However, in real-world scenarios, individuals can have diversified preferences on the same topic based on their different experiences, educational backgrounds, religions, and cultures (Leonardelli et al., 2021). Even for the same person, his or her expected model answer to a particular question can vary depending on different scenarios (Cheng et al., 2023a). The annotation disagreement, which is caused by different annotators or the same annotator in different scenarios Bai et al. (2022), will significantly hinder the effectiveness of alignment methods Davani et al. (2022); Wan et al. (2023); He et al. (2024).
@@
In our exploration, we found the over-rewarding phenomenon, that is, the vanilla RMs tend to output extreme rewards on samples, which damages the RMs and LLM alignment. To enhance the efficiency of leveraging the diversified preference datasets, inspired by multi-objective optimization methods Sener and Koltun (2018); Zeng et al. (2023c), we regard RMs as a shared reward additionally with a customized reward drift. The shared reward represents the shared preferences across datasets (or general human preferences) and the reward drift contains individual or domain-specific preference information (Cheng et al., 2023a). Then, we introduce a Multi-Objective Reward training scheme (MORE) to capture the shared (general) preference information, which adopts a novel reweight techniques to minimize the mean gradient of enlarging reward drifts. With MORE, RMs can capture a broader range of preferences and mitigate the impact of reward drifts. The main contributions of this paper are:
@@ 4.3 Training Scheme: MORE
where we adjust to minimize the partial gradient of enlarging reward drifts. The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023c). We provide the details of our efficient implementation in the Appendix B. Furthermore, shares the same magnitude of vanilla loss function in expectation over the whole training dataset, as justified in Appendix B.
@@ 5 Experiments on Reward Modeling
In this section, we present our experiments and quantitative analyses on reward modeling. The open-source code and data are available at https://github.com/dunzeng/MORE.
@@ 5.1 Reward Modeling on Diversified Preference Datasets
Increasing the size of LLMs, reward model training on mixed diversified preference datasets can maintain reward accuracy. For instance, when Pythia-1.4B is used as the RM base model, the reward accuracy is lower than the Top accuracy achieved through single preference training on all preferences. Then, when LLaMa2-7B is used as the base model, the reward accuracy on the Oasst1, Webgpt, and Summarise test sets surpasses the top accuracy achieved through single training. Additionally, the degradation of reward accuracy on the Helpful and Harmless datasets is mitigated. Therefore, the performance of RMs typically is proportional to the size of base models (Gao et al., 2023). Moreover, we find the accuracy of RM is low, revealing the preference conflicts across RM.
@@
We provide additional reward modeling results of LLaMa2-13B in the Appendix. Compared with the results of LLaMa2-7B, the reward accuracy of LLaMa2-13B is not significantly better. This is because the capacity of these 7B and 13B models is sufficient for fitting the datasets used. Notably, the ECE of the 13B model is marginally improved in the MultiTask setting, and the ECE gap between RM and RM is narrowed. Thus, a larger reward model can mitigate the negative impacts of mixed diverse preferences.
Reward Model Perplexity (PPL) GPT4 Evaluation (%) Base Model Scheme Acc(%) ECE Helpful Harmless Win Tie Lose - - - - 15.48 12.71 - - - Pythia-1.4B MultiTask 64.79 0.0177 15.30 8.22 44 22 34 MORE 64.32 0.0109 12.68 8.42 Pythia-2.8B MultiTask 66.61 0.0145 16.76 8.42 45 21 34 MORE 65.87 0.0078 13.14 10.29 LLaMa2-7B MultiTask 72.40 0.0284 16.93 8.69 45 23 32 MORE 72.32 0.0143 11.97 9.96 Table 1: The RJS alignment performance with different RMs. The first line is the performance of the Alpaca base model. The results show that ECE further reflects the ability of RMs when the reward accuracy is close.
@@ 6 Experiments on LLM Alignment
We finetune the Alpaca model on the good response from H&H training datasets, and the finetuned model is marked by Alpaca-SFT. Then, we conduct the RJS alignment experiments with LLaMa2-7B RMs from Figure 6. In Figure 7, we compare each alignment result of Alpaca-RJS models with the same Alpaca-STF model via GPT evaluation (the tie rates are around 15%). The results show that the RMs with lower ECE values work better for RJS alignments, emphasizing the importance of calibration evaluation.
@@ 7 Additional Discussions
Connections with fine-grained reward and multi-dimensional reward
Existing research, including fine-grained reward (Wu et al., 2024) and multi-dimensional reward (Lou et al., 2024), has increasingly focused on the importance of reward diversity. These studies categorize the utilization of diverse reward signals into two main strategies: first, integrating multi-dimensional preferences into a single dimension for aligning large language models (LLMs); and second, decoupling preferences across dimensions (reward models) to align LLMs on each dimension. This body of work underscores the necessity of addressing diverse preferences. In contrast, our primary contribution lies in the empirical analysis of diversified preferences within a single reward model, alongside our novel evaluation of expected calibration error (ECE) on reward models. Additionally, since both fine-grained and multi-dimensional reward methodologies yield scalar rewards, our findings regarding ECE are relevant for assessing these approaches.
@@
This paper reveals two main suggestions for future reward model training works. First, Evaluate RMs with reward accuracy and calibration error. Reward accuracy is insufficient to evaluate the ability of RMs due to model capacity and data quality. Our work suggests that the community additionally focuses on the calibration performance of RMs. Besides, Increasing the diversity of preference data samples can ensure the robustness of the reward modeling process. Due to the preference information being typically noisy, learning reward information from mixed diversified datasets can be beneficial.
@@
The MORE can enhance preference modeling pre-trained (PMP) paradigm (Askell et al., 2021) as it captures the shared preference information. This facilitates its use in federated learning scenarios (McMahan et al., 2017; Zeng et al., 2023b, a), where the data distributions are highly heterogeneous across participants. Moreover, the RM can be easily finetuned to specific preferences (Cheng et al., 2023a). This flexibility allows for the adaptation of our approach to various applications.
@@ Appendix A Related Work
Multi-objective Optimization (MOO) (Gunantara, 2018) is a branch of methods addressing learning problems involving multiple conflicting objectives. In real-world scenarios, it commonly encounters situations where multiple objectives need to be considered simultaneously, often with trade-offs between them. In the practice of machine learning, most MOO methods Sener and Koltun (2018); Zeng et al. (2023c) apply linear scalarization Barrett and Narayanan (2008) to merge multiple objectives into one, and then automatically adjust the objective coefficients to balance the conflicts among different tasks.
@@ Appendix B Detailed Discussions about MORE
We use adaptive weighting methods to reduce the reward drift across preferences and adjust the reward modeling process in the data batch-wise. The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023c). However, the computing cost of solving it is proportional to the size of parameters . Since the size of is in the billions, we only utilize gradients on the reward head from each preference to avoid expensive computation cost. In detail, we obtain the hidden states before the reward head and compute the gradient of the reward head solely with data . Collecting the reward head gradient from diversified preferences, the is computed by: